# Quantitative imaging of nonlinear spin-wave propagation using diamond quantum sensors

### v1

@@ Abstract
Spin waves propagating in magnetic materials exhibit nonlinear behavior due to the competition between excitation and relaxation at large amplitudes, providing an attractive platform for exploring nonlinear wave dynamics. In particular, spin waves with a non-zero wavenumber that carry momentum undergo nonlinear relaxation and experience wavenumber modulation in the nonlinear regime. We image the nonlinear spin-wave propagation in two yttrium iron garnet thin films with distinct spin-wave decay rates using a wide-field quantum diamond microscope. We obtain quantitative distributions of spin-wave amplitude and phase in varying excitation microwave strength. As a result, we observe a threshold in the spin-wave amplitude beyond which nonlinear effects become evident and confirm that this threshold is consistent with theoretical predictions based on four-magnon scattering for both samples. Moreover, as the amplitude of the spin waves increases, modulation of the wavenumber is observed across the field of view. We attribute this modulation primarily to the spin waves generated by multi-magnon scattering. Our quantitative measurements provide a pathway for visualizing nonlinear spin-wave dynamics and are crucial for deepening our understanding of the underlying mechanisms.
March 30, 2025
@@ I Introduction
Spin waves, collective excitations of ordered magnets, possess intrinsic nonlinearity [1, 2]. This nonlinearity emerges from the competition between spin-wave excitation and relaxation and becomes prominent when the spin-wave amplitude exceeds a certain threshold. The nonlinear dynamics of the uniform mode (i.e., ferromagnetic resonance, FMR) have been comprehensively studied from the 1950s to the present. In these studies, various nonlinear processes have been explored, including parametric excitation [3, 4, 5], Bose-Einstein condensation of magnons [6], and the first and second Suhl instabilities [1]. Recently, significant interest has focused on the dynamics of spin waves with non-zero wavenumbers due to their rich physics and potential applications in information devices [7]. The nonlinear behavior of these spin waves has been investigated both theoretically and experimentally [8, 9, 10, 11, 12, 13], in which the initially excited coherent spin waves are scattered by multi-magnon processes, thereby magnons with various wavenumbers and frequencies are generated. In addition, when the spin-wave amplitude becomes large, an effective reduction in the static magnetization leads to a frequency shift of the spin waves [2, 13, 14].
These nonlinear dynamics of the spin waves are primarily investigated using a vector network analyzer (VNA) [12], magneto-optical effects [15], and, especially in recent years, Brillouin light scattering (BLS) [8, 9, 10, 13]. BLS enables time-, space-, and frequency-resolved detection of spin-wave propagation, making it a valuable technique for visualizing these dynamics [16]. However, the signal obtained via BLS is the intensity of the scattered light, which arises from Raman scattering processes. While the intensity is proportional to the magnon intensity, there is no direct theoretical relationship connecting it to the spin-wave amplitude. Although Ref. [13] estimates the precession angle at which nonlinearity emerges for the spin waves in a magnetic waveguide by calibrating the spin-wave amplitude based on variation in the wavenumber, quantitative comparisons with theoretical predictions have not been performed, and thus the validity of the obtained amplitude has not been verified.
@@
Although some previous research utilizing NV centers has observed nonlinear spin-wave dynamics, such as generation of incoherent spin waves at the resonant frequencies of the NV spins when exciting the FMR mode [23, 24] and frequency conversion of coherent spin waves by spin-wave mixing through four-magnon scattering [25], no studies to date have quantitatively examined the nonlinear dynamics of propagating spin waves with non-zero wavenumbers themselves.
Figure 1: (a) Schematic of the experiment. (b) The dispersion relation of the surface spin waves for YIG-A at external magnetic field $B_{0}=25.17\,\mathrm{mT}$. The horizontal dashed line corresponds to the resonance frequency of the NV spin. (c) Two-dimensional image of the microwave amplitude at an external magnetic field $B_{0}=25.17\,\mathrm{mT}$ and input microwave power $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ for YIG-A. (d) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for the data in (c).
In this paper, we present results of imaging the amplitude and phase of surface spin waves in two yttrium iron garnet (YIG) thin films with distinct characteristics using a wide-field quantum diamond microscope (QDM) while systematically varying the excitation microwave amplitude. First, we show that the spin waves in the first YIG thin film grown by sputtering exhibit significant decay beyond a characteristic threshold. We confirm this threshold is consistent with theoretical predictions based on four-magnon scattering. In addition, as the amplitude of the spin waves increases, we observe modulation of the wavenumber within the field of view, and our quantitative analysis reveals that this modulation largely derives from the spin waves that are generated by multi-magnon scattering. Finally, to validate our findings, we show that the second YIG thin film, which is grown by liquid phase epitaxy (LPE), and exhibiting a smaller relaxation rate compared to the first YIG film, exhibits a distinct spin-wave amplitude threshold, also in agreement with theoretical expectations. This achievement demonstrates the usefulness of quantitative QDM investigations and contributes to a systematic understanding of the nonlinear spin waves.
This paper is organized as follows. We describe the experimental setup and principle in Sec. II. In Sec. III.1, we show spin-wave imaging in the linear regime for the first YIG film. We present the nonlinear behavior when the amplitude of the spin waves increases in Sec. III.2, and extract the spin-wave amplitude in Sec. III.3. Then, we introduce the theoretical formulation to discuss the threshold between the linear and nonlinear regime and apply it to the experimental findings in Sec. III.4. We address the modulation of the wavenumber in Sec. III.5. Sec. III.6 is devoted to comparing with the second YIG film. Finally, we conclude the paper with perspectives in Sec. IV.
@@ II Methods
The experimental setup is depicted in Fig. 1(a), where we employ QDM for the measurement system. In this setup, a green laser (wavelength: 515nm) irradiates an ensemble of NV centers, and their red fluorescence is collected using an sCMOS camera [26]. The magnification of the microscope is 100x, and the field of view is approximately $100\,\mathrm{\mu m}$ square.
@@
We use two YIG thin films with different growth methods as our magnetic materials. The first film, “YIG-A”, is epitaxially grown on (111)-oriented gadolinium gallium garnet (GGG) substrate using RF sputtering in the group of the co-authors of this paper [28] and has a thickness of $d=54\,\mathrm{nm}$. The second film, “YIG-B”, is epitaxially grown on a (111)-oriented GGG substrate using LPE and is commercially obtained from Matesy GmbH, with a thickness of $d=109\,\mathrm{nm}$. For the YIG-A, the measurement setup is identical to that described in Ref. [26] except for the configuration of the microwave circuits.
Striplines with a width of approximately $20\,\mathrm{\mu m}$ are deposited on the film surface for both YIG samples. Microwaves are input to the stripline on the left side to excite the spin waves [Fig. 1(a)]. We define the edge of the $+x$ side of this stripline as the origin of the $x$-axis ($x=0$). The striplines on YIG-A consist of Ti/Cu/Au ($10\,\mathrm{nm}/190\,\mathrm{nm}/9\,\mathrm{nm}$), and those on YIG-B consist of Ti/Cu/Au ($10\,\mathrm{nm}/97\,\mathrm{nm}/9\,\mathrm{nm}$). Microwaves are first amplified from the signal generator through an amplifier and are then fed into the coplanar waveguide; from there, they are delivered to the stripline via gold wires attached using a wire bonder. We define the input microwave power $P_{\mathrm{mw}}$ as the power supplied to the coplanar waveguide. Consequently, due to differences in losses, the actual microwave power reaching the stripline may differ between YIG-A and YIG-B.
The diamond chips are attached to the YIG films by varnish, with the surface containing the NV layer facing downward. The distance $z_{\mathrm{d}}$ between the NV layer and the surface of the magnetic film is estimated by evaluating the distribution of the stray magnetic field from the stripline on the surface of the YIG film under current application [26]. For YIG-A and YIG-B, the distance $z_{d}$ is estimated to be $(878\pm 20)\,\mathrm{nm}$ and $z_{d}=(860\pm 18)\,\mathrm{nm}$, respectively (see Supplemental Materials for details).
A static magnetic field is applied in the film plane along the stripline ($y$-axis) using a Helmholtz coil. In this configuration, surface spin waves propagating perpendicular to the magnetic field are observed [29].
@@ III.1 Spin-wave imaging in the linear regime
where $\gamma_{e}$ is the gyromagnetic ratio of an electron spin, $\mu_{0}$ is the permeability of vacuum, $M_{\mathrm{eff}}\,(=169.6\,\mathrm{mT})$ is the effective magnetization [26], and $d=54\,\mathrm{nm}$ is the thickness of the YIG film. In this magnetic field, the resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2165.8\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.4\,\mathrm{rad/\mu m}$.
Figure 1(c) shows the two-dimensional distribution of the microwave field amplitude at an input microwave power of $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ under $B_{0}=25.17\,\mathrm{mT}$. In the analysis, signals from each $520\,\mathrm{nm}$ in the $x$- and $y$-directions are integrated and treated as a single pixel. The microwave amplitude oscillates periodically in the $x$-direction due to the interference between the microwave generated by the spin waves propagating in the $x$-direction and the reference microwave with a spatially uniform phase across the entire field of view [21, 22, 26]. In addition to the propagation in the $x$-direction, an increase in the microwave amplitude is observed at a specific diagonal angle. This phenomenon, known as “caustics,” arises from the anisotropy of the spin-wave dispersion relation [34, 21, 26].
Figure 1(d) displays the one-dimensional microwave amplitude distribution along the $x$-direction $B_{\mathrm{mw}}(x)$. Here, signals are integrated over a width of $13\,\mathrm{\mu m}$ from $y=31.87\,\mathrm{\mu m}$ to $y=44.87\,\mathrm{\mu m}$ in the $y$-direction. The leftmost $x$-coordinate of the experimental data is $x=0.65\,\mathrm{\mu m}$. The microwave amplitude at the position of the NV centers results from the interference between the microwave field generated by the spin waves and that from the stripline and is expressed as [21, 22, 26]
@@
where $B_{\mathrm{strip}}(x)$ is the microwave amplitude distribution from the stripline, $B_{\mathrm{sw}}(x)$ is that from the spin waves, $k_{x}(x)$ is spatial distribution of the $x$-component of the wavenumber and $\theta_{0}$ is the phase of the spin waves at the left edge of the field of view. When $B_{\mathrm{sw}}$ ($B_{\mathrm{strip}}$) is sufficiently larger than $B_{\mathrm{strip}}$ ($B_{\mathrm{sw}}$), $B_{\mathrm{sw}(\mathrm{strip})}\gg B_{\mathrm{strip}(\mathrm{sw})}$, the above equation can be approximated as
@@
For YIG-A, under low input microwave power, the condition $B_{\mathrm{sw}}(x)>B_{\mathrm{mw}}(x)$ holds in the entire field of view (see Supplemental Materials for details). Consequently, the mean line (the line passing through the center of the oscillations) of $B_{\mathrm{mw}}(x)$ shown in Fig. 1(d) corresponds to $B_{\mathrm{sw}}(x)$, while the oscillatory component corresponds to $B_{\mathrm{strip}}(x)$. Based on the decay of this mean line, we can estimate that the decay length of the spin waves is approximately $50\,\mathrm{\mu m}$.
@@ III.2 From linear to nonlinear regime
We examine the case when the input power is varied in 10 steps from $5\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$. Figure 2(a) displays $B_{\mathrm{mw}}(x)$ at five selected power values (see Supplemental Materials for all data). The triangular markers with error bars represent the experimental data, while the solid black lines indicate the fitting results discussed later. The integrated region along the $y$-direction is the same as in Fig. 1(d). From $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to about $P_{\mathrm{mw}}=19\,\mathrm{dBm}$, the shape of the amplitude distribution remains unchanged; however, for input power above this range, significant modulation in the distribution is observed.
Figure 2: (a) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained at five different input microwave powers from $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$ for YIG-A. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data between large input microwave power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$) and small input microwave power ($P_{\mathrm{mw}}=11\,\mathrm{dBm}$) for YIG-A. (c) Microwave amplitude distribution and (d) Wavenumber distribution from the spin waves for each microwave input power calculated from the fitting results for YIG-A.
Figure 2(b) compares $B_{\mathrm{mw}}(x)$ measured at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=25\,\mathrm{dBm}$. The first noticeable point is that the shapes of the mean lines of $B_{\mathrm{mw}}(x)$ differ between the two power levels. This variation arises because $B_{\mathrm{strip}}(x)$ increases linearly with the input microwave power, whereas $B_{\mathrm{sw}}(x)$ exhibits a nonlinear response. According to Eq. (3), at low input powers, the mean line primarily reflects $B_{\mathrm{sw}}(x)$. In contrast, at high input powers, the mean line is dominated by the contribution from the stripline (see Supplemental Materials for details). Furthermore, in the region of smaller $x$-coordinate [see the inset of Fig. 2(b)], the spacing between the peaks of the microwave amplitude oscillations is noticeably shorter at $P_{\mathrm{mw}}=25\,\mathrm{dBm}$ compared to $11\,\mathrm{dBm}$. This increase in the wavenumber is likely due to a modulation of the dispersion relation, which is caused by a reduction in the static magnetization along the $y$-direction due to the increase in the spin-wave amplitude [13].
@@
where $x_{0}$ denotes the coordinate of the left edge of the field of view ($x_{0}=0.65\,\mathrm{\mu m}$ in this measurement). For $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$, we adopt a stretched exponential function. The function for the spin waves (stripline) is characterized by the amplitude at $x=x_{0}$, $B^{0}_{\mathrm{sw}}=B_{\mathrm{sw}}(x_{0})$ ($B^{0}_{\mathrm{strip}}=B_{\mathrm{strip}}(x_{0})$), the decay length $l_{\mathrm{sw}}$ ($l_{\mathrm{strip}}$), and the stretch factor $p_{\mathrm{sw}}$ ($p_{\mathrm{strip}}$). In the linear region, $B_{\mathrm{sw}}(x)$ generally decays exponentially; however, in the nonlinear region, rapid decay of the amplitude near the stripline has been reported [13]. Therefore, we introduce a stretch factor into the exponential function to capture this behavior (see Supplemental Materials for details). For the wavenumber distribution, we adopt a function consisting of a stretched exponential function corresponding to the spatial modulation of the wavenumber and the constant term $k_{0}$. The stretched exponential function is characterized by the wavenumber modulation at $x=x_{0}$, $k_{1}$, the decay length $l_{k}$, and the stretch factor $p_{\mathrm{k}}$.
@@
When performing the fitting, we must note that Eq. (2) is symmetric with respect to $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$. Starting from random initial values makes it impossible to distinguish between these two contributions. To resolve this ambiguity, we use the microwave amplitude distribution obtained under the same microwave input power but with the direction of the external magnetic field reversed as a reference. When the magnetic field is reversed, the amplitude of the spin waves propagating in the $+x$-direction decreases due to the reduction of the spin-wave excitation efficiency of the stripline [21, 35]. Also, the microwave from the spin waves propagating in the $+x$-direction in this configuration does not induce transitions between the NV spins’ $m_{s}=0$ and $m_{s}=-1$ states due to its polarization [21]. Thus, we assume that the reference microwave amplitude distribution is primarily due to the stripline. Accordingly, we first fit the reference distribution and then use the obtained fitting parameters as the initial parameters for $B_{\mathrm{strip}}(x)$ in Eq. (4) (see Supplemental Materials for details).
@@
The resulting fittings are shown in Fig. 2(a) as solid black lines. For each input microwave power, the fitted curve reproduces the experimental data well. In Fig. 2(c), we present the estimated microwave amplitude distribution from the spin waves $B_{\mathrm{sw}}(x)$ obtained from these fitting results. Here, we define the left end of the plot at $x=0.91\,\mathrm{\mu m}$, corresponding to the region where the amplitude variation from the stripline becomes relatively small. For input powers ranging from $P_{\mathrm{mw}}=5\,\mathrm{dBm}$ to approximately $19\,\mathrm{dBm}$, $B_{\mathrm{sw}}(x)$ increases with increasing input power and decays gradually over a similar decay length within the field of view. However, at microwave powers above this range, it does not continue to increase and instead decays rapidly near the stripline. In Fig. 2(d), we show the wavenumber distribution $k_{x}(x)$ estimated by fitting for each input microwave power. As the input power increases, the wavenumber also increases. While no significant spatial variation is observed within the field of view at lower input powers, noticeable spatial variations emerge from around $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. At the maximum power of $25\,\mathrm{dBm}$, the wavenumber varies by more than 20% across the field of view.

### v2

@@ Abstract
Spin waves propagating in magnetic materials exhibit nonlinear behavior at large amplitudes due to the competition between excitation and relaxation, providing an attractive platform for exploring nonlinear wave dynamics. In particular, spin waves with a non-zero wavenumber that carry momentum undergo nonlinear relaxation and experience wavenumber modulation in the nonlinear regime. This nonlinearity has been observed experimentally—for example, in S. R. Lake et al., Phys. Rev. Appl. 17, 034010 (2022)— but a quantitative comparison with theory has not yet been carried out. Here, We image nonlinear spin-wave propagation in two yttrium iron garnet thin films with distinct spin-wave decay rates using a wide-field quantum diamond microscope. We obtain quantitative distributions of spin-wave amplitude and phase as a function of the excitation microwave strength. As a result, we observe a threshold in the spin-wave amplitude beyond which nonlinear effects become evident and confirm that this threshold is consistent with theoretical predictions based on four-magnon scattering for both samples. Moreover, as the amplitude of the spin waves increases, we observe modulation of the wavenumber across the field of view. We attribute this modulation primarily to a reduction in the saturation magnetization caused by incoherent spin waves generated by multi-magnon scattering. Our quantitative measurements provide a pathway for visualizing nonlinear spin-wave dynamics and are crucial for deepening our understanding of the underlying mechanisms.
December 11, 2025
@@ I Introduction
Spin waves, collective excitations of ordered magnets, possess an intrinsic nonlinearity [1, 2]. This nonlinearity arises from the competition between spin-wave excitation and relaxation and becomes prominent when the spin-wave amplitude exceeds a certain threshold. The nonlinear dynamics of the uniform mode (i.e., ferromagnetic resonance, FMR) have been comprehensively studied since the 1950s. In these studies, various nonlinear processes have been explored, including parametric excitation [3, 4, 5], Bose-Einstein condensation of magnons [6], and the first and second Suhl instabilities [1]. Recently, significant interest has focused on the dynamics of spin waves with non-zero wavenumbers due to their rich physics and potential applications in information devices [7]. The nonlinear behavior of these spin waves has been investigated both theoretically and experimentally [8, 9, 10, 11, 12, 13], in which the initially excited coherent spin waves are scattered by multi-magnon processes, thereby magnons with various wavenumbers and frequencies are generated. In addition, when the spin-wave amplitude becomes large, an effective reduction in the static magnetization leads to a frequency shift of the spin waves [2, 13, 14].
These nonlinear dynamics of the spin waves are primarily investigated using a vector network analyzer (VNA) [12], magneto-optical effects [15], and, especially in recent years, Brillouin light scattering (BLS) [8, 9, 10, 13]. BLS enables time-, space-, and frequency-resolved detection of spin-wave propagation, making it a valuable technique for visualizing these dynamics [16]. However, the signal obtained via BLS is the intensity of the scattered light arising from Raman scattering processes. While this intensity is proportional to the magnon intensity, there is no direct theoretical relationship that links it to the spin-wave amplitude. Although Ref. [13] systematically investigates the nonlinearity of surface spin waves propagating on a magnetic waveguide, discusses the underlying mechanism, and estimates the precession angle at which nonlinearity emerges by calibrating the spin-wave amplitude from the variation in wavenumber, quantitative comparisons with theoretical predictions have not been carried out, and thus the validity of the obtained amplitude has not yet been verified.
@@
Although some previous studies employing NV centers has observed nonlinear spin-wave dynamics, such as the generation of incoherent spin waves at the resonant frequencies of the NV spins when exciting the FMR mode [23, 24] and the frequency conversion of coherent spin waves by spin-wave mixing through four-magnon scattering [25], no study to date has quantitatively examined the nonlinear dynamics of propagating spin waves with non-zero wavenumbers themselves.
Figure 1: (a) Schematic of the experiment. (b) Dispersion relation of the surface spin waves for YIG-A at an external magnetic field $B_{0}=25.17\,\mathrm{mT}$. The horizontal dashed line corresponds to the resonance frequency of the NV spin. (c) Two-dimensional image of the microwave amplitude at an external magnetic field $B_{0}=25.17\,\mathrm{mT}$ and input microwave power $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ for YIG-A. (d) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for the data in (c).
In this paper, we present results on imaging the amplitude and phase of surface spin waves in two yttrium iron garnet (YIG) thin films with distinct characteristics using a wide-field quantum diamond microscope (QDM), while systematically varying the excitation microwave amplitude. First, we show that the spin waves in the first YIG thin film, grown by sputtering, exhibit a significant decay beyond a characteristic amplitude threshold. We confirm that this threshold is consistent with theoretical predictions based on four-magnon scattering. In addition, as the amplitude of the spin waves increases, we observe modulation of the wavenumber within the field of view, and our quantitative analysis reveals that this modulation largely originates from incoherent spin waves generated by multi-magnon scattering. Finally, to validate our findings, we show that the second YIG thin film, which is grown by liquid phase epitaxy (LPE) and characterized by a smaller relaxation rate than the first YIG film, exhibits a distinct spin-wave amplitude threshold, also in agreement with theoretical expectations. This achievement demonstrates the usefulness of quantitative QDM investigations for probing nonlinear spin-wave dynamics and contributes to a systematic understanding of nonlinear spin waves.
This paper is organized as follows. We describe the experimental setup and principle in Sec. II. In Sec. III.1, we show spin-wave imaging in the linear regime for the first YIG film. We then present the nonlinear behavior that emerges as the spin-wave amplitude is increased in Sec. III.2, and quantitatively extract the spin-wave amplitude in Sec. III.3. Then, we introduce the theoretical formulation to discuss the threshold between the linear and nonlinear regimes and apply it to our experimental findings in Sec. III.4. We address the modulation of the wavenumber in Sec. III.5. Sec. III.6 is devoted to comparing with the second YIG film. Finally, we conclude the paper with a summary and outlook in Sec. IV.
@@ II Methods
The experimental setup is depicted in Fig. 1(a), where we employ a QDM as the measurement system. In this setup, a green laser (wavelength: $515\,\mathrm{nm}$) irradiates an ensemble of NV centers, and their red fluorescence is collected using an sCMOS camera [26]. The magnification of the microscope is $100\times$, and the field of view is approximately $100\,\mathrm{\mu m}$ square.
@@
We use two YIG thin films with different growth methods as our magnetic materials. The first film, “YIG-A”, is epitaxially grown on (111)-oriented gadolinium gallium garnet (GGG) substrate using RF sputtering in the group of the co-authors of this paper [28], and has a thickness of $d=54\,\mathrm{nm}$. The second film, “YIG-B”, is epitaxially grown on a (111)-oriented GGG substrate by LPE and is commercially obtained from Matesy GmbH, with a thickness of $d=109\,\mathrm{nm}$. For YIG-A, the measurement setup is identical to that described in Ref. [26] except for the configuration of the microwave circuits.
Striplines with a width of approximately $20\,\mathrm{\mu m}$ and a length of $2\,\mathrm{mm}$, with $500\,\mathrm{\mu m}\times 500\,\mathrm{\mu m}$ pads for wiring at both ends, are deposited on the film surface for both YIG samples. Microwaves are input to the stripline on the left side to excite spin waves [Fig. 1(a)]. The striplines on YIG-A consist of Ti/Cu/Au ($10\,\mathrm{nm}/190\,\mathrm{nm}/9\,\mathrm{nm}$), and those on YIG-B consist of Ti/Cu/Au ($10\,\mathrm{nm}/97\,\mathrm{nm}/9\,\mathrm{nm}$). Microwaves are first amplified from the signal generator through an amplifier and are then fed into the coplanar waveguide; from there, they are delivered to the stripline by wire bonding gold wires to the pad. The other side is also connected to the coplanar waveguide and terminated with $50\,\Omega$. We define the input microwave power $P_{\mathrm{mw}}$ as the power supplied to the coplanar waveguide. Consequently, due to differences in losses, the actual microwave power reaching the stripline may differ between YIG-A and YIG-B.
The diamond chips are attached to the YIG films using varnish, with the surface containing the NV layer facing downward. The distance $z_{d}$ between the NV layer and the surface of the magnetic film is estimated by evaluating the distribution of the stray magnetic field from the stripline on the surface of the YIG film under current application [29, 26]. For YIG-A and YIG-B, the distance $z_{d}$ is estimated to be $(878\pm 20)\,\mathrm{nm}$ and $z_{d}=(860\pm 18)\,\mathrm{nm}$, respectively (see Supplemental Materials for details [30]).
A static magnetic field is applied in the film plane along the stripline ($y$-axis) using a Helmholtz coil. In this configuration, surface spin waves propagating perpendicular to the magnetic field are excited [31].
@@ III.1 Spin-wave imaging in the linear regime
where $\gamma_{e}$ is the gyromagnetic ratio of an electron spin, $\mu_{0}$ is the permeability of vacuum, $M_{\mathrm{eff}}\,(=169.6\,\mathrm{mT})$ is the effective magnetization [26], and $d=54\,\mathrm{nm}$ is the thickness of the YIG film. Under this magnetic field, the resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2165.8\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.4\,\mathrm{rad/\mu m}$.
Figure 1(c) shows the two-dimensional distribution of the microwave field amplitude at an input microwave power of $P_{\mathrm{mw}}=11\,\mathrm{dBm}$. In the analysis, signals from each $520\,\mathrm{nm}$ in the $x$- and $y$-directions are integrated and treated as a single pixel. The microwave amplitude oscillates periodically in the $x$-direction due to the interference between the microwave field generated by the spin waves propagating in the $x$-direction and the reference microwave with a spatially uniform phase across the entire field of view [21, 22, 26]. In addition to the propagation in the $x$-direction, an increase in the microwave amplitude is observed at a specific diagonal angle. This phenomenon, known as “caustics,” arises from the anisotropy of the spin-wave dispersion relation [36, 21, 26].
Figure 1(d) displays the one-dimensional microwave amplitude distribution along the $x$-direction $B_{\mathrm{mw}}(x)$. Here, signals are integrated over a width of $13\,\mathrm{\mu m}$ from $y=31.87\,\mathrm{\mu m}$ to $y=44.87\,\mathrm{\mu m}$ in the $y$-direction. The leftmost $x$-coordinate of the experimental data is $x=0.65\,\mathrm{\mu m}$. The microwave amplitude at the position of the NV centers results from the interference between the microwave field generated by the spin waves and that from the stripline, and can be expressed as [21, 22, 26]
@@
where $B_{\mathrm{strip}}(x)$ is the microwave amplitude distribution from the stripline, $B_{\mathrm{sw}}(x)$ is that from the spin waves, $k_{x}(x)$ is the spatial distribution of the $x$-component of the wavenumber and $\theta_{0}$ is the phase of the spin waves at the left edge of the field of view. When $B_{\mathrm{sw}}$ ($B_{\mathrm{strip}}$) is sufficiently larger than $B_{\mathrm{strip}}$ ($B_{\mathrm{sw}}$), $B_{\mathrm{sw}(\mathrm{strip})}\gg B_{\mathrm{strip}(\mathrm{sw})}$, the above equation can be approximated as
@@
For YIG-A, under low input microwave power, the condition $B_{\mathrm{sw}}(x)>B_{\mathrm{mw}}(x)$ holds in the entire field of view (see Supplemental Materials for details [30]). Consequently, the mean line (the line passing through the center of the oscillations) of $B_{\mathrm{mw}}(x)$ shown in Fig. 1(d) corresponds to $B_{\mathrm{sw}}(x)$, while the oscillatory component corresponds to $B_{\mathrm{strip}}(x)$. From the decay of this mean line, we can estimate that the decay length of the spin waves is approximately $50\,\mathrm{\mu m}$.
@@ III.2 From linear to nonlinear regime
We examine the case where the input power is varied in 10 steps from $5\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$. Figure 2(a) displays $B_{\mathrm{mw}}(x)$ at five selected power values (see Supplemental Materials for the full dataset [30]). The triangular markers with error bars represent the experimental data, while the solid black lines indicate the fitting results discussed later. The integrated region along the $y$-direction is the same as in Fig. 1(d). From $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to about $P_{\mathrm{mw}}=17\,\mathrm{dBm}$, the shape of the amplitude distribution remains unchanged; however, for input power above this range, significant modulation of the distribution is observed.
Figure 2: (a) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained at five different input microwave powers from $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$ for YIG-A. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data between large input microwave power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$) and small input microwave power ($P_{\mathrm{mw}}=11\,\mathrm{dBm}$) for YIG-A. (c) Microwave amplitude and (d) Wavenumber distribution from the spin waves for each input microwave power calculated from the fitting results for YIG-A.
Figure 2(b) compares $B_{\mathrm{mw}}(x)$ measured at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=25\,\mathrm{dBm}$. The first notable feature is that the shapes of the mean lines of $B_{\mathrm{mw}}(x)$ differ between the two power levels. This variation arises because $B_{\mathrm{strip}}(x)$ increases linearly with the input microwave power, whereas $B_{\mathrm{sw}}(x)$ exhibits a nonlinear response. According to Eq. (3), at low input powers, the mean line primarily reflects $B_{\mathrm{sw}}(x)$. In contrast, at high input powers, the mean line is dominated by the contribution from the stripline (see Supplemental Materials for details [30]). Furthermore, in the region of smaller $x$-coordinate [see the inset of Fig. 2(b)], the spacing between the peaks of the microwave amplitude oscillations is noticeably shorter at $P_{\mathrm{mw}}=25\,\mathrm{dBm}$ compared to $11\,\mathrm{dBm}$. This increase in the wavenumber is likely due to a modulation of the dispersion relation, caused by a reduction in the static magnetization along the $y$-direction resulting from the increase in the spin-wave amplitude [13].
@@
where $x_{0}$ denotes the coordinate of the left edge of the field of view ($x_{0}=0.65\,\mathrm{\mu m}$ in this measurement). For $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$, we adopt stretched exponential functions. The function for the spin waves (stripline) is characterized by the amplitude at $x=x_{0}$, $B^{0}_{\mathrm{sw}}=B_{\mathrm{sw}}(x_{0})$ ($B^{0}_{\mathrm{strip}}=B_{\mathrm{strip}}(x_{0})$), the decay length $l_{\mathrm{sw}}$ ($l_{\mathrm{strip}}$), and the stretch factor $p_{\mathrm{sw}}$ ($p_{\mathrm{strip}}$). In the linear regime, $B_{\mathrm{sw}}(x)$ generally decays exponentially; however, in the nonlinear regime, rapid decay of the amplitude near the stripline has been reported [13]. Therefore, we introduce a stretch factor into the exponential function to capture this behavior (see Supplemental Materials for details [30]). For the wavenumber distribution, we adopt a function consisting of a stretched exponential function corresponding to the spatial modulation of the wavenumber and a constant term $k_{0}$. The stretched exponential function is characterized by the wavenumber modulation at $x=x_{0}$, $k_{1}$, the decay length $l_{k}$, and the stretch factor $p_{\mathrm{k}}$.
@@
When performing the fitting, we note that Eq. (2) is symmetric with respect to $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$. Starting from random initial values makes it impossible to distinguish between these two contributions. To resolve this ambiguity, we use the microwave amplitude distribution obtained under the same input microwave power but with the direction of the external magnetic field reversed as a reference. When the magnetic field is reversed, the amplitude of the spin waves propagating in the +x+x-direction decreases due to a reduction of the spin-wave excitation efficiency of the stripline [21, 37]. Moreover, the microwave field from the spin waves propagating in the +x+x-direction in this configuration does not induce transitions between the NV spins’ $m_{s}=0$ and $m_{s}=-1$ states because of its polarization [21, 38]. Thus, we assume that the reference microwave amplitude distribution arises primarily from the stripline. Accordingly, we first fit the reference distribution and then use the obtained fitting parameters as the initial values for $B_{\mathrm{strip}}(x)$ in Eq. (4) (see Supplemental Materials for details [30]).
@@
The fitted curves are shown in Fig. 2(a) as solid black lines. For each input microwave power, the fitted curve reproduces the experimental data well. In Fig. 2(c), we present the estimated microwave amplitude distribution from the spin waves $B_{\mathrm{sw}}(x)$ obtained from these fitting results. Here, we define the left end of the plot at $x=0.91\,\mathrm{\mu m}$, corresponding to the region where the amplitude variation from the stripline becomes relatively small. For input powers ranging from $P_{\mathrm{mw}}=5\,\mathrm{dBm}$ to approximately $19\,\mathrm{dBm}$, $B_{\mathrm{sw}}(x)$ increases with increasing input power and decays gradually over a similar decay length within the field of view. However, at microwave powers above this range, it no longer increases and instead decays rapidly near the stripline. In Fig. 2(d), we show the wavenumber distribution $k_{x}(x)$ estimated by fitting for each input microwave power. As the input power increases, the wavenumber also increases. While no significant spatial variation is observed within the field of view at lower input powers, noticeable spatial variations emerge from around $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. At the maximum power of $25\,\mathrm{dBm}$, the wavenumber varies by more than 20% across the field of view.